OpenStudy (anonymous):

7 over y+2 times 4 over y-2 equals 11 over y squared - 4

OpenStudy (anonymous):

please write using symbol......

OpenStudy (anonymous):

\[7 \div (y+2) \times 4 \div (y-2) = 11 \div (y^{2} - 4)\] notice \[(y+2)(y-2) = y^{2} - 4\] but your equation is incorrect because you end up with 28 = 11 which is false